Circular No. 7098

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1550–564 B. A. Harmon, M. H. Finger, M. L. McCollough, S. N. Zhang, W. S.

Paciesas, and C. A. Wilson, Marshall Space Flight Center, NASA, report, for the Compton Gamma Ray Observatory/BATSE Team: “The x-ray transient XTE J1550–564 (IAUC 7008) was detected in BATSE data beginning on Jan. 21. As of Jan. 23, the one-day averaged flux was 300 mCrab (± 20 percent, 20–100 keV), with a hard spectrum (power-law photon-number index −2.3 ± 0.2). This is a reflare in hard x-rays, following the primary outburst in 1998 Sept.–Oct. (IAUC 7010). The source has been active in soft x-rays (2–12 keV) since ∼ Dec. 18, according to observations by the Rossi X-ray Timing Explorer ASM.”

GRB 990123 W. Offutt, Cloudcroft, NM, obtained two 600-s and two 900-s expo-

sures of the optical candidate for GRB 990123, using a 0.60-m f/7 Ritchey-Chrétien reflector (+ CCD + Bessell R-band filter). The mean position from the four exposures is α = 15^{h}25^{m}30^{s}.35, δ = +44^{o}45′59′′.2 (equinox 2000.0; twelve USNO SA2.0 stars as reference). The following photometry (mean residual 0.25 mag) is based on the reference star located ∼ 11′′.4 west of the target object (R = 19.1): Jan. 24.4400 UT, 19.8; 24.4567, 19.8; 24.4666, 20.0; 24.4781, 20.3.

B. A. Skiff, Lowell Observatory, writes, in reference to IAUC 7096: “In the absence of a proper calibration sequence, optical observers of GRB 990123 have been making photometric zero-point adjustments using USNO A2.0 magnitudes. It is perhaps worth noting that these data do not in general correspond to either the Johnson B or Cousins R systems, and that they often possess substantial systematic errors in scale, in the zeropoint, and as a function of star color. For the GSPC sequence near the center of the relevant plate, plus a few other stars from the Minnesota APS

group (http://isis.spa.umn.edu/Docs/get photom.html), it turns out

that the A2.0 ‘mb’ values are close to standard B; ‘mr’ is offset ∼ +0.3 mag from Cousins R (A2.0 being too faint). More typically, both ‘mb’ and ‘mr’ are ∼ 0.6 mag too bright compared to B and R. My large file of faint photometric reference stars can be used to make local corrections to A2.0 or oth-

er catalogues of faint stars: ftp://ftp.lowell.edu/pub/bas/starcats/
loneos.stds.”

1999 January 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green